e Purchased supplies Tor $105,000 on account. 4. Prepare a trial balance at July 31. Credit BEARINGS & BRAKES CORPORATION TriBearings & Brakes Corporation (B&B) was incorporated as a private company. The company’s accounts included the following at June 30: Accounts Payable $ 88,000 Buildings 650,000 Cash 105,000 Common Stock 320,000 Equipment 178,000 Land 519,000 Notes Payable (long-term) 8,000 Retained Earnings 1,041,000 Supplies 5,000 During the month of July, the company had the following activities: Issued 4,400 shares of common stock for $440,000 cash. Borrowed $135,000 cash from a local bank, payable in four years. Bought a building for $196,000; paid $81,000 in cash and signed a three-year note for the balance. Paid cash for equipment that cost $105,000. Purchased supplies for $105,000 on account.
Answer #1
BEARINGS & BRAKES CORPORATION
Trial Balance
For the Month Ended July 31
Debit Credit
Accounts Payable $       193,000
Buildings $         846,000
Cash $         494,000
Common Stock $       760,000
Equipment $         283,000
Land $         519,000
Notes Payable (Long Term) $       258,000
Retained Earnings $    1,041,000
Supplies $         110,000
Total $     2,252,000 $    2,252,000

Here is the formula version

BEARINGS & BRAKES CORPORATION
Trial Balance
For the Month Ended July 31
Debit Credit
Accounts Payable =88000+105000
Buildings =650000+196000
Cash =105000+440000+135000-81000-105000
Common Stock =320000+440000
Equipment =178000+105000
Land 519000
Notes Payable (Long Term) =135000+8000+115000
Retained Earnings 1041000
Supplies =5000+105000
Total =SUM(D8:D18) =SUM(E8:E18)